# Longitudinal Abuse and Sentiment Analysis of Hollywood Movie Dialogues using LLMs

### v2

@@ 4.4 Case studies
We visualised some dialogues from four classic movies in Tables 3, 4, 5, and 6, corresponding to the time points of the extreme values of sentiment polarity and abusive content count in Figures 14, 15, 16, and 17.
@@
In Table 3, The Departed demonstrates a wide range of emotional tones. At 00:27:32,684 to 00:27:34,777, ”Yeah, that’s right, fancy. Are you a statie?” expresses medium-level positive emotion, possibly indicating curiosity or mild amusement. However, from 01:32:36,031 to 01:32:40,434, the tone shifts dramatically with ”You pressure me to fear for my life and I will put a fucking bullet in your head…”, which conveys a high level of negative emotion, likely reflecting intense anger or threat. This intensity continues at 01:33:05,393 to 01:33:07,691 with ”One of your guys is going to pop you,” which also shows high negative emotion, suggesting violence or hostility. The sentiment remains negative from 01:33:10,365 to 01:33:12,663, where ”As for running drugs, what the fuck are you doing?” conveys frustration and strong disapproval. At 01:33:12,834 to 01:33:16,429, ”You don’t need the money or the pain in the ass. And they will catch you.” expresses medium-level negative emotion, indicating resignation or warning. Finally, at 01:33:23,211 to 01:33:26,476, ”Tell you the truth, I don’t need pussy anymore, either” reveals a medium-level negative tone, possibly reflecting a sense of emotional detachment or bitterness. Overall, the emotional tone shifts sharply from medium positive to high negative, conveying strong hostility and frustration throughout the dialogue.
@@
Annie Hall’s sentiment shifts range from positive to negative in Table 4. In the beginning, from 00:25:34,241 to 00:25:39,036, the phrase ”You play well” reflects a low level of positive emotion. Then, from 00:34:44,415 to 00:34:46,416, ”Give me a kiss” conveys a medium level of positive emotion. As the time progresses, the tone turns negative, with ”Sometime you make me so mad!” from 00:44:05,142 to 00:44:09,270 and the following lines showing low-level negative emotion. This is followed by more negative sentiment in 00:52:36,820 to 00:52:40,197, criticising adult education with medium-level negativity. At 00:55:30,118 to 00:55:33,329, the statement about the Stones concert expresses low negative emotion. A shift occurs at 01:25:05,308 to 01:25:07,309, where ”You know how wonderful you are” introduces a high level of positive emotion. In the next few timestamps, the tone turns negative again, especially in 01:28:28,761 to 01:28:31,013 with references to ”death,” and in 01:29:13,097 to 01:29:15,432, with ”difficult in life” showing medium-level negativity. Overall, this dialogue showcases significant emotional fluctuations from positive to negative and back to positive, highlighting the complexity of interactions and the evolution of emotions.
@@
In Table 5, the emotional tone of The Silence of the Lambs fluctuates between positive and negative. In 00:30:20,985 to 00:30:24,029, ”Dr. Chilton does enjoy his petty torments” expresses a low level of positive emotion, possibly hinting at amusement or satisfaction. This contrasts with the high level of negative emotion at 01:10:58,337 to 01:11:01,089, where the statement ”They were slaughtering the spring lambs?” carries a strong negative sentiment, likely reflecting shock or horror. At 01:32:04,519 to 01:32:07,103, the mention of ”Starling, we know who he is and where he is” shows a low level of positive emotion, suggesting a sense of control or certainty. The emotional tone turns sharply negative again at 01:35:33,477 to 01:35:35,311 with the phrase ”I’d fuck me so hard,” reflecting high negative emotion, likely in a disturbed or unsettling context. Finally, at 01:45:32,034 to 01:45:35,077, ”How the fuck should I know? Just get me out of here!” expresses a high level of negative emotion, indicating frustration and desperation. Overall, the dialogue demonstrates a strong contrast between low positive and high negative emotions throughout these moments.
@@
In Table 6, from 00:10:44,140 to 00:10:48,020, ”The enemy of one the enemy of all is” expresses a medium level of negative emotion, likely conveying a sense of conflict or tension. This is followed by another negative statement at 00:10:59,260 to 00:11:01,700, where ”They kicked me out of. Ring-a-ring o’Roses” also carries medium negative emotion, possibly suggesting rejection or frustration. The tone then shifts to positive starting at 00:45:48,800 to 00:45:51,200, with ”Expecting great things, from all accounts” showing a low level of positive emotion, hinting at optimism or hope. At 00:46:00,080 to 00:46:02,600, ”And good luck for this afternoon. Thank you.” expresses medium-level positive emotion, offering encouragement and good wishes. Finally, at 01:25:59,280 to 01:26:03,760, ”Then let’s hope that’s wise enough to give you room for manoeuvre” expresses a low level of positive emotion, suggesting cautious optimism or hope for flexibility. Overall, the emotional tone moves from medium negative to positive, reflecting shifts in expectations and interactions.
@@ 4.4.1 Take some examples to conduct sentiment analysis and abusive detection for the entire movie
Figure 14 shows the change in sentiment polarity over time for ”The Departed”, which is an action and drama movie. The sentiment polarity fluctuates considerably throughout the movie’s timeline and peaks around the 30th minute. Furthermore, the abusive words are frequent during certain periods of the film, especially between 65-105 minutes and 135-145 minutes. These periods may correspond to high-intensity dialogues or conflict scenes. A combination of the two graphs suggests a possible correlation between significant decreases in emotional polarity and a high frequency of abusive content or behaviour.
@@
Figure 15 presents sentiment polarity and abusive content for the movie ”Annie Hall”. The emotional polarity of the movie fluctuates within a generally positive range, mostly between -0.05 and 0.20. Around the 90th minute, the polarity briefly drops into a negative zone but quickly rebounds, reflecting an emotional low point followed by recovery. This trend aligns with the movie’s overall optimistic tone, likely tied to its comedic and lighthearted theme. We find that the abusive content is concentrated and sporadic, which shows that only a certain part of the film contains relatively concentrated violent or offensive language content. We find no significant change in sentiment polarity during the period of high incidence of violent content (about 40 to 60 minutes), but there was a small positive fluctuation in emotional polarity during this period. The sharp decline in emotional polarity at the 90th minute is not accompanied by the detection of abusive content, suggesting that other non-violent emotional factors, this period may have character conflict and inner struggle in the story.
@@
Figure 16 presents sentiment polarity and abusive content detection for the ”Silence of the Lambs”. We observe that the sentiment polarity undergoes significant fluctuations at several points, with more pronounced positive peaks, especially around periods 30, 65, and 90, suggesting that these moments likely contain more positive emotional content. We can observe abusive content that appears in different periods of the movie. Among them, the time period between 95 and 105 minutes has the most violent content, indicating that this is the film’s climax and could be the key violent scene appearing. The violent content also increased during periods 65 and 85, but the intensity of violence was relatively low compared to the peak of 95-105. This may indicate that the film tension is elevated at these moments, but not at its highest point.
@@
Figure 17 presents sentiment polarity for ”Chariots of Fire” which peaks at the 240th and 260th minutes capturing moments of heightened emotional intensity to represent emotional climaxes within the film. The abusive content detection indicates shows that there is almost no violence in the film and only a very small amount appears in approximately 150, 215, and 275 minutes. This validates that most of the film revolves around the sportsmanship and inner struggles of the characters, rather than outright violence or conflict.
@@ 5 Discussion
Our research shows that sentiment polarity analysis underscores the nuanced emotional landscape of films, clearly indicating that movies have become more emotionally complex over the years. The average sentiment scores reveal a trend towards more balanced emotional expressions, with a slight decline in positive sentiment polarity in recent decades. This could be attributed to the increasing depiction of complex and often darker themes in contemporary cinema. The sentiment co-occurrence analysis revealed intriguing patterns in how emotions interplay within movie dialogues. High co-occurrence frequencies of emotions like joking and annoyance, or sadness and anger, highlight the complex emotional dynamics filmmakers utilise to engage audiences. The rarity of co-occurrences between emotions like thankfulness and pessimism further illustrates the deliberate emotional structuring in cinematic storytelling. In related studies, Palomo et al. [87] used deep learning technology to conduct sentiment analysis on IMDB movie reviews and revealed the changes in emotional tendencies in movie reviews of different periods. They reported that from the 1950s to the 1980s, emotional expression underwent significant changes, and since the 1980s, emotional expression has become more complex, especially under the influence of social events such as the global financial crisis, climate change and pandemics. The impact of negative emotions has become more pronounced given feat from catastrophic natural events, poor governance and politics. However, our study of movie subtitles takes a different approach where we evaluated the content of the movie rather than how the society views them.
@@
Regarding detection of abuse, our study overlaps with Demarty et al. [88] who provided a study on multimodal violence detection by combining video, audio and text subtitle analysis. They reported that the accuracy of violence detection can be significantly improved by integrating visual, audio and language features. Information from the different modes complements each other, helping to capture more complex features of violent scenes. However, our approach pays more attention to the combination of text and context. In our case, we observed that there  has been fluctuations in levels of violence in the two major categories (Oscar-nominated and blockbusters) and the level of abuse for Oscar-movies took over blockbusters since 2010 (Figure 11). This could be due to over-commercialisation, rise of social media with mobile phones, and competition with OTT [89, 90] and other streaming platforms including Instagram Stories, YouTube Reels, and Tiktok. Sharma et al. [91] highlighted the rise of abuse and violence in relation to women, caste and religion through Indian web series in OTT platforms. Due to the easy accessibility of OTT platforms, cinema release movies have been facing challenges in attracting viewers to the cinema, since the movies are later released in OTT. The normalisation of abuse and hate-speech, particularly in OTT platforms could be one of the reasons the traditional movies have an increase in abusive content over the last decade, and due to their popularity they have been making it to Oscar awards.
@@
Although our study provides a comprehensive overview of the emotional and abusive content in Hollywood films, there are several limitations. Firstly, the movie data is limited with subtitles of Oscar nominations and the top yearly blockbuster movies. Therefore, they do not cover all types of films and cultural backgrounds. Oscar films tend to have a specific artistic and cultural bent, while the top ten films at the box office may be more commercial and entertaining. As a result, these data may not reflect niche films’ emotional and abusive content characteristics. In addition, we selected only films with English subtitles, excluding other languages, contributing to our findings’ limitations. Secondly, as a comprehensive art form, the audience is not only affected by dialogue but also by a variety of factors including vision. We only analysed the film’s subtitles, ignoring visual information and description of the scenes where no dialogue takes place. The description of violence in the film is usually conveyed through visual effects, including action scenes, blood effects, and violent action scenes. Finally, there are limitations to our choice of fine-tuning LLM datasets. The data we selected for fine-tuning abusive detection came from messages on the online community Reddit from January 2012 to June 2015, which is not comprehensive. The meaning and use of abusive content have changed significantly over time. Some words may be considered highly insulting in one era, but may have lost their original strong insult in another. In addition, the abusive content dataset used in the study is largely based on contemporary language norms, which may not accurately reflect abusive words and their contexts from different eras in history.
@@
The use of abusive and vulgar language in cinema movies and OTT platforms has become a topic of controversy and debate in recent years [89]. Unlike traditional television networks, OTT platforms are not subject to the same regulations, allowing them greater freedom in producing and distributing content. Fortunately, most OTT platforms have age ratings and content warnings to help viewers make informed decisions about the content they consume. In response to concerns about using vulgar language on OTT platforms, government regulators have introduced new rules and guidelines [92, 93]. Developing rules and guidelines to regulate content on OTT platforms is an issue that requires urgent attention [89] with more and more people using OTT platforms. We need to establish a strict content vetting mechanism to ensure that all content uploaded to the platform is vetted. The vetting should include legal compliance, morality and ethics, violence and pornographic content, etc., to safeguard the legality and compliance of the content.
@@
In future research, sentiment analysis and violence detection can use multimodal data analysis [94, 95], that is, not only text data but also visual and audio data. More comprehensive emotional information can be obtained by analysing visual and auditory elements (such as expressions, tone of voice, etc.) in movies. In addition, the research can further refine the classification of emotion and can be extended to films in different cultural and regional backgrounds to compare the expression of emotion and the use of abusive content in different cultures. This will help to understand the cultural influence on the film’s content. At the same time, developing a real-time analysis system is also a necessary research direction, which can conduct instant emotional and abusive content analysis of newly released movies, and provide immediate review and recommendation to subscribers. Finally, through the knowledge of psychology, sociology and other fields, in-depth research can be done on the impact of film content on audience psychology and behaviour, especially the impact of abusive and violent content on teenagers.
6 Conclusion

### v3

@@ 4.4 Case studies
We presented some dialogues from four classic movies in Tables 5, 6, 7, and 8, corresponding to the time points of the extreme values of sentiment polarity and abusive content count in Figures 14, 15, 16, and 17.
@@
In Table 5, The Departed demonstrates a wide range of emotional tones. Time frame 00:27:32,684 to 00:27:34,777, ”Yeah, that’s right, fancy. Are you a statie?” expresses medium-level positive emotion, possibly indicating curiosity or mild amusement. However, in time frame 01:32:36,031 to 01:32:40,434, the tone shifts dramatically with ”You pressure me to fear for my life and I will put a fucking bullet in your head…”, which conveys a high level of negative emotion, likely reflecting intense anger or threat. This intensity continues in the time frame 01:33:05,393 to 01:33:07,691 with ”One of your guys is going to pop you,” which also shows high negative emotion, suggesting violence or hostility. The sentiment remains negative from 01:33:10,365 to 01:33:12,663, where ”As for running drugs, what the fuck are you doing?” conveys frustration and strong disapproval. Time frame 01:33:12,834 to 01:33:16,429, ”You don’t need the money or the pain in the ass. And they will catch you.” expresses medium-level negative emotion, indicating resignation or warning. Finally, time frame 01:33:23,211 to 01:33:26,476, ”Tell you the truth, I don’t need pussy anymore, either” reveals a medium-level negative tone, possibly reflecting a sense of emotional detachment or bitterness. Overall, the emotional tone shifts sharply from medium positive to high negative, conveying strong hostility and frustration throughout the dialogue.
@@
Annie Hall’s sentiment shifts range from positive to negative in Table 6. In the beginning, from time frame 00:25:34,241 - 00:25:39,036, the phrase ”You play well” reflects a low level of positive emotion. Then, from 00:34:44,415 to 00:34:46,416, ”Give me a kiss” conveys a medium level of positive emotion. As the time progresses, the tone turns negative, with ”Sometime you make me so mad!” from 00:44:05,142 to 00:44:09,270, and the following lines showing low-level negative emotion. This is followed by more negative sentiment in 00:52:36,820 to 00:52:40,197, criticising adult education with medium-level negativity. At 00:55:30,118 to 00:55:33,329, the statement about the Stones concert expresses low negative emotion. A shift occurs at 01:25:05,308 to 01:25:07,309, where ”You know how wonderful you are” introduces a high level of positive emotion. In the next few timestamps, the tone turns negative again, especially in 01:28:28,761 to 01:28:31,013 with references to ”death,” and in 01:29:13,097 to 01:29:15,432, with ”difficult in life” showing medium-level negativity. Overall, this dialogue showcases significant emotional fluctuations from positive to negative and back to positive, highlighting the complexity of interactions and the evolution of emotions.
@@
In Table 7, the emotional tone of The Silence of the Lambs fluctuates between positive and negative. In time frame 00:30:20,985 to 00:30:24,029, ”Dr. Chilton does enjoy his petty torments” expresses a low level of positive emotion, possibly hinting at amusement or satisfaction. This contrasts with the high level of negative emotion at time frame 01:10:58,337 - 01:11:01,089, where the statement ”They were slaughtering the spring lambs?” carries a strong negative sentiment, likely reflecting shock or horror. At time frame 01:32:04,519 - 01:32:07,103, the mention of ”Starling, we know who he is and where he is” shows a low level of positive emotion, suggesting a sense of control or certainty. The emotional tone turns sharply negative again at time frame 01:35:33,477 - 01:35:35,311 with the phrase ”I’d fuck me so hard,” reflecting high negative emotion, likely in a disturbed or unsettling context. Finally, at time frame 01:45:32,034 - 01:45:35,077, ”How the fuck should I know? Just get me out of here!” expresses a high level of negative emotion, indicating frustration and desperation. Overall, the dialogue demonstrates a strong contrast between low positive and high negative emotions throughout these moments.
@@
In Table 8, the time frame 00:10:44,140 - 00:10:48,020, ”The enemy of one the enemy of all is” expresses a medium level of negative emotion, likely conveying a sense of conflict or tension. This is followed by another negative statement at time frame 00:10:59,260 - 00:11:01,700, where ”They kicked me out of. Ring-a-ring o’Roses” also carries medium negative emotion, possibly suggesting rejection or frustration. The tone then shifts to positive, starting at time frame 00:45:48,800 - 00:45:51,200, with ”Expecting great things, from all accounts” showing a low level of positive emotion, hinting at optimism or hope. At time frame 00:46:00,080 -00:46:02,600, ”And good luck for this afternoon. Thank you.” expresses medium-level positive emotion, offering encouragement and good wishes. Finally, at time frame 01:25:59,280 - 01:26:03,760, ”Then let’s hope that’s wise enough to give you room for manoeuvre” expresses a low level of positive emotion, suggesting cautious optimism or hope for flexibility. Overall, the emotional tone moves from medium negative to positive, reflecting shifts in expectations and interactions.
@@ 4.4.1 Take some examples to conduct sentiment analysis and abusive detection for the entire movie
Figure 14 shows the change in sentiment polarity over time for The Departed, which is an action and drama movie. The sentiment polarity fluctuates considerably throughout the movie’s timeline and peaks around the 30th minute. Furthermore, the abusive words are frequent during certain periods of the film, especially between 65-105 minutes and 135-145 minutes. These periods may correspond to high-intensity dialogues or conflict scenes. A combination of the two graphs suggests a possible correlation between significant decreases in emotional polarity and a high frequency of abusive content or behaviour.
@@
Figure 15 presents sentiment polarity and abusive content for the movie Annie Hall. The emotional polarity of the movie fluctuates within a generally positive range, mostly between -0.05 and 0.20. Around the 90th minute, the polarity briefly drops into a negative zone but quickly rebounds, reflecting an emotional low point followed by recovery. This trend aligns with the movie’s overall optimistic tone, likely tied to its comedic and lighthearted theme. The abusive content is concentrated and sporadic, which shows that only a certain part of the film contains relatively concentrated violent or offensive language content. No significant change in sentiment polarity during the period of high incidence of violent content (about 40 to 60 minutes), but there was a small positive fluctuation in emotional polarity during this period. The sharp decline in emotional polarity at the 90th minute is not accompanied by the detection of abusive content, suggesting that other non-violent emotional factors, this period may have character conflict and inner struggle in the story.
@@
Figure 16 presents sentiment polarity and abusive content detection for The Silence of the Lambs. The sentiment polarity undergoes significant fluctuations at several points, with more pronounced positive peaks, especially around periods 30, 65, and 90, suggesting that these moments likely contain more positive emotional content. Abusive content that appears in different periods of the movie. Among them, the time period between 95 and 105 minutes has the most violent content, indicating that this is the film’s climax and could be the key violent scene. The violent content also increased during periods 65 and 85, but the intensity of violence was relatively low compared to the peak of 95-105. This may indicate that the film’s tension is elevated at these moments, but not at its highest point.
@@
Figure 17 presents sentiment polarity for Chariots of Fire which peaks at the 240th and 260th minutes capturing moments of heightened emotional intensity to represent emotional climaxes within the film. The abusive content detection indicates shows that there is almost no violence in the film, and only a very small amount appears in approximately 150, 215, and 275 minutes. This validates that most of the film revolves around the sportsmanship and inner struggles of the characters, rather than outright violence or conflict.
@@ 5 Discussion
Our research shows that sentiment polarity analysis underscores the nuanced emotional landscape of films, clearly indicating that movies have become more emotionally complex over the years. The average sentiment scores reveal a trend towards more balanced emotional expressions, with a slight decline in positive sentiment polarity in recent decades. This could be attributed to the increasing depiction of complex and often darker themes in contemporary cinema. The sentiment co-occurrence analysis revealed intriguing patterns in how emotions interplay within movie dialogues. High co-occurrence frequencies of emotions like joking and annoyance, or sadness and anger, highlight the complex emotional dynamics filmmakers utilise to engage audiences. The rarity of co-occurrences between emotions like thankfulness and pessimism further illustrates the deliberate emotional structuring in cinematic storytelling. In related studies, Palomo et al. (Palomo et al.,, 2024) used deep learning technology to conduct sentiment analysis on IMDb movie reviews and revealed the changes in emotional tendencies in movie reviews of different periods. They reported that from the 1950s to the 1980s, emotional expression underwent significant changes, and since the 1980s, emotional expression has become more complex, especially under the influence of social events such as the global financial crisis, climate change and pandemics. The impact of negative emotions has become more pronounced, given the rise of films featuring catastrophic natural events, poor governance and politics.
@@
Regarding detection of abuse, our study overlaps with Demarty et al. (Demarty et al.,, 2014) who provided a study on multimodal violence detection by combining video, audio and text subtitle analysis. They reported that the accuracy of violence detection can be significantly improved by integrating visual, audio and language features. A multimodal detection framework (Wu et al.,, 2020) from 2020 showed that combining audio and visual cues under weak supervision significantly improved violence recognition performance. Information from the different modes complements each other, helping to capture more complex features of violent scenes. However, our approach pays more attention to the combination of text and context. There have been fluctuations in levels of violence in the two major categories (Oscar-nominated and blockbusters), and the level of abuse for Oscar movies took over blockbusters since 2010 (Figure 11). This could be due to over-commercialisation, the rise of social media with mobile phones, and competition with OTT (Dhiman,, 2023; Simlote et al.,, 2024) and other streaming platforms, including Instagram Stories, YouTube Reels, and TikTok. Sharma et al. (Sharma,, 2021) highlighted the rise of abuse and violence in relation to women, caste and religion through Indian web series on OTT platforms. Due to the easy accessibility of OTT platforms, cinema-released movies have been facing challenges in attracting viewers to the cinema, as the movies are often released on OTT platforms shortly after their cinema release. The normalisation of abuse and hate speech, particularly in OTT platforms, could be one of the reasons traditional movies have seen an increase in abusive content over the last decade, and due to their popularity, they have been making it to the Oscar awards.
Our findings offer several practical implications for stakeholders in the film and media industry. First, film producers and streaming platforms can leverage the observed trends in sentiment polarity and abusive content to refine their content moderation, audience targeting, and storytelling strategies. For example, the increase in abusive language in recent decades highlights the importance of adopting more sophisticated content filtering and age-rating systems on OTT platforms. Second, regulators may consider how the meaning of abusive language evolves when designing classification guidelines, ensuring that content ratings remain relevant to contemporary social norms. Finally, the observed cross-genre and award/commercial differences suggest that marketing strategies should be genre-sensitive, leveraging insights into emotional dynamics to engage audiences more effectively.
@@
Although our study provides a comprehensive overview of the emotional and abusive content in Hollywood films, there are several limitations. Firstly, the movie data is limited to subtitles of Oscar nominations and the top yearly blockbuster movies. Therefore, they do not cover all types of films and cultural backgrounds. Oscar films tend to have a specific artistic and cultural bent, while the top ten films at the box office may be more commercial and entertaining. As a result, these data may not reflect niche films’ emotional and abusive content characteristics. In addition, we only selected films with English subtitles, excluding other languages, contributing to our findings’ limitations. Secondly, as a comprehensive art form, the audience is not only affected by dialogue but also by a variety of factors, including vision. We only analysed the film’s subtitles, ignoring visual information and description of the scenes where no dialogue takes place. The description of violence in the film is usually conveyed through visual effects, including action scenes, blood effects, and violent action scenes. Finally, there are limitations regarding the datasets used for fine-tuning language models. The data to fine-tune abusive detection came from messages on the Reddit online community from January 2012 to June 2015, which is not comprehensive. The meaning and use of abusive content have changed significantly over time. Some words may be considered highly insulting in one era, but may have lost their original strong insult in another. In addition, the abusive content dataset used in the study is largely based on contemporary language norms, which may not accurately reflect abusive words and their contexts from different eras in history.
@@
The use of abusive and vulgar language in cinema movies and OTT platforms has become a topic of controversy and debate in recent years (Dhiman,, 2023). Unlike traditional television networks, OTT platforms are not subject to the same regulations, allowing them greater freedom in producing and distributing content. Fortunately, most OTT platforms have age ratings and content warnings to help viewers make informed decisions about the content they consume. In response to concerns about using vulgar language on OTT platforms, government regulators have introduced new rules and guidelines (Baccarne et al.,, 2013; Baldry et al.,, 2014). Developing rules and guidelines to regulate content on OTT platforms is an issue that requires urgent attention (Dhiman,, 2023) with more and more people using OTT platforms. A strict content vetting mechanism needs to be established to ensure that all content uploaded to the platform is vetted. The vetting should include legal compliance, morality and ethics, violence and pornographic content, etc., to safeguard the legality and compliance of the content.
Another limitation arises from potential biases in both the subtitle corpus and the fine-tuning datasets. Subtitles from OpenSubtitles may differ in accuracy, with potential omissions, simplifications, or translator-specific interpretations, which could affect the fidelity of dialogue analysis. Moreover, our fine-tuning corpora (SenWave and RAL-E) are rooted in English-language social media, limiting the generalizability of findings to other linguistic and cultural settings. As a result, while our study offers meaningful insights into Hollywood English films, caution must be applied when extrapolating results to non-English or culturally diverse contexts. Future research could incorporate multilingual subtitle corpora and culturally diverse datasets to improve robustness and cross-cultural generalizability.
@@
In future research, sentiment analysis and violence detection can use multimodal data analysis (Ngiam et al.,, 2011; Tavakoli et al.,, 2025), that is, not only text data but also visual and audio data. More comprehensive emotional information can be obtained by analysing visual and auditory elements (such as expressions, tone of voice, etc.) in movies. In addition, the research can further refine the classification of emotion and can be extended to films in different cultural and regional backgrounds to compare the expression of emotion and the use of abusive content in different cultures. This will help to understand the cultural influence on the film’s content. At the same time, developing a real-time analysis system is also a necessary research direction, which can conduct instant emotional and abusive content analysis of newly released movies, and provide immediate review and recommendation to subscribers. Finally, through the knowledge of psychology, sociology and other fields, in-depth research can be done on the impact of film content on audience psychology and behaviour, especially the impact of abusive and violent content on teenagers.
6 Conclusions